# Multilayered Intelligent Reflecting Surface for Long-Range Underwater Acoustic Communication

### v1

@@ VI-A1 Experimental settings
The reflectors consist of two PZT-4 disks separated by a polyethylene terephthalate (PET) film. The sound wave propagation speed in the disk is approximately 3900 m/s; each PZT disk measures $\phi$ 49.2 mm $\times$ 2.7 mm and resonates near 44 kHz. The matching network and load states of the reflectors are controlled by the circuit board shown in Fig. 3.
VI-A2 Preprocessing
@@
$\left\{\begin{array}[]{lll}\vspace{0.2cm}\!\!r_{opop}(t)=\mathcal{A}(t)+\!\displaystyle\sum^{M}_{k=1}\!\left(B_{1}^{op}e^{j\phi_{1}^{op}}\!+\!B_{2}^{op}e^{j\phi_{2}^{op}}\right)\!s(t\!-\!\tau_{k}),\\ \vspace{0.13cm}\!\!r_{shsh}(t)=\mathcal{A}(t)+\!\displaystyle\sum^{M}_{k=1}\!\left(B_{1}^{sh}e^{j\phi_{1}^{sh}}\!\!\!+\!B_{2}^{sh}e^{j\phi_{2}^{sh}}\right)\!s(t\!-\!\tau_{k}),\\ \end{array}\right.$ (11)
where $r_{opop}(t)$ and $r_{shsh}(t)$ represent the received signals when the two PZT disks are in open-circuit and short-circuit status, respectively. The term $s(t)$ is the original transmission signal and $\mathcal{A}(t)\!\!=\!\!\!\sum^{N}_{i=0}\!A_{i}e^{j\psi_{i}}s(t\!-\!\tau_{i})$, where $A_{i}$, $\psi_{i}$, and $\tau_{i}$ denote the amplitude, phase, and propagation delay of the channel coefficients associated with the direct wave ($i\!=\!0$) and the $i$th multipath signal ($i\!>\!0$) that do not interact with the reflector. The parameters $B_{1}^{op}$ and $\phi_{1}^{op}$ specify the amplitude and phase, respectively, of the wave reflected by the first PZT layer under open-circuit status, while $B_{2}^{op}$ and $\phi_{2}^{op}$ represent those for the second PZT layer.
@@
$\mathcal{A}(t)=\frac{r_{opop}(t)+r_{shsh}(t)}{2}.$ (12)
From (11), it can be observed that $\mathcal{A}(t)$ does not depend on the reflector’s load status in a stable acoustic environment. Therefore, the waves reflected by the reflector for any load status can be extracted from the received signal using two reference signals. For instance, if the loads of the first and second layers of the PZT stack are $\text{C}_{\text{0.6}}$ and $\text{R}_{\text{L}}$, respectively, the received signal is denoted by $r_{C_{0.6}R_{L}}(t)$, while the signal reflected by the PZT stack is given by
$\mathcal{B}_{\,C_{0.6}R_{L}}(t)\!=\!\displaystyle\sum^{M}_{k=1}\!\left(B^{C_{0.6}}_{1}e^{j\phi^{C_{0.6}}_{1}}\!+B^{R_{L}}_{2}e^{j\phi^{R_{L}}_{2}}\right)\!s(t-\tau_{k}).$ (13)
Based on (12), $\mathcal{B}_{\,C_{0.6}R_{L}}(t)$ can be calculated as follows:
$\begin{array}[]{lll}\mathcal{B}_{\,C_{0.6}R_{L}}(t)\!\!\!\!&=&\!\!\!r_{C_{0.6}R_{L}}(t)-\mathcal{A}(t)\\ \!\!\!\!&=&\!\!\!r_{C_{0.6}R_{L}}(t)-\displaystyle\frac{r_{opop}(t)+r_{shsh}(t)}{2}.\end{array}$ (14)
The expression in (14) is used to extract the signal reflected by the reflector in a multipath environment from experimental measurements. For performance evaluation, the theoretical result of the reflected wave, $\mathcal{B}^{*}_{\,C_{0.6}R_{L}}(t)$, is also needed. According to the circuit design presented in Fig. 3, under ideal conditions,
$\left\{\begin{array}[]{lll}\vspace{0.2cm}\!\!\!B^{C_{0.6}}_{1}e^{j\phi^{C_{0.6}}_{1}}\!\!\!\!\!\!\!&=&\!\!\!0.6\,e^{-j\frac{\pi}{2}},\\ \vspace{0.13cm}\!\!\!B^{R_{L}}_{2}e^{j\phi^{R_{L}}_{2}}\!\!\!\!\!\!\!\!&=&\!\!\!\displaystyle\frac{R_{L}-Z_{0}}{R_{L}+Z_{0}},\\ \end{array}\right.$ (15)
@@
$\mathcal{B}^{*}_{\,C_{0.6}R_{L}}(t)\!=\!\left[0.6e^{-j\frac{\pi}{2}}\!+\!\left(\displaystyle\frac{R_{L}-Z_{0}}{R_{L}+Z_{0}}\right)\right]\displaystyle\sum^{M}_{k=1}\!s(t\!-\!\tau_{k}).$ (16)
A comparison of $\mathcal{B}_{\,C_{0.6}R_{L}}(t)$ with $\mathcal{B}^{*}_{\,C_{0.6}R_{L}}(t)$ enables evaluating the accuracy of the amplitude and phase of the reflected wave. For convenience, we first normalize $\mathcal{B}_{\,C_{0.6}R_{L}}(t)$ and $\mathcal{B}^{*}_{\,C_{0.6}R_{L}}(t)$ through $\mathcal{B}_{\,opop}(t)$. Here, $\mathcal{B}_{\,opop}(t)$ represents the signal reflected by the PZT stack when both layers are in open-circuit status, i.e., $\mathcal{B}_{\,opop}(t)\!=\!\sum^{M}_{k=1}\!\left(B^{op}_{1}e^{j\phi^{op}_{1}}\!\!\!+\!B^{op}_{2}e^{j\phi^{op}_{2}}\right)\!s(t-\tau_{k})$, which can be obtained experimentally as follows:
$\mathcal{B}_{\,opop}(t)=\displaystyle\frac{r_{opop}(t)-r_{shsh}(t)}{2}=2\displaystyle\sum^{M}_{k=1}\!s(t-\tau_{k}).$ (17)
According to the (16) and (17), the normalized $\mathcal{B}^{*}_{\,C_{0.6}R_{L}}(t)$ can be expressed as follows:
$\frac{\mathcal{B}^{*}_{\,C_{0.6}R_{L}}(t)}{\mathcal{B}_{\,opop}(t)}=0.3\,e^{-j\frac{\pi}{2}}\!+\!\displaystyle\frac{R_{L}-Z_{0}}{2\left(R_{L}+Z_{0}\right)}$ (18)
Finally, when the load status of the two layers are $\text{C}_{\text{0.6}}$ and $\text{R}_{\text{L}}$, the accuracy of the reflected signal generated by the multilayer reflector in a multipath environment can be estimated by the difference between $\mathcal{B}^{*}_{\,C_{0.6}R_{L}}(t)/\mathcal{B}_{\,opop}(t)$, calculated by (18), with $\mathcal{B}_{\,C_{0.6}R_{L}}(t)/\mathcal{B}_{\,opop}(t)$, obtained experimentally via (14). This methodology can be applied to analyze the amplitude and phase of the reflection signals for arbitrary load status in the multilayered acoustic reflector.
VI-A3 Experimental results
@@
Fig. 10 (b) illustrates the waveform of $\mathcal{A}(t)$, which includes the direct wave and the multiple signals that do not interact with the reflector. As described in (12), $\mathcal{A}(t)$ is computed as half the sum of the two reference waves shown in Fig. 10 (a). In Fig. 10 (c), we display the waves reflected by the PZT stack under two different load conditions, where $\mathcal{B}_{\,C_{0.9}R_{\text{2k}}}(t)$ and $\mathcal{B}_{\,opop}(t)$ are obtained from the two reference waveforms through the expressions introduced in (14) and (17), respectively.
To demonstrate the difference between the two reflected signals, Fig. 10 (d) provides a partially enlarged view of the area within the yellow dotted frame in Fig. 10 (c). In this figure, let $\Delta\phi_{\,C_{0.9}R_{\text{2k}}}$ denote the phase difference between $\mathcal{B}_{\,opop}(t)$ and $\mathcal{B}_{\,C_{0.9}R_{\text{2k}}}(t)$. Experimental results indicate that the average $\Delta\phi_{\,C_{0.9}R_{\text{2k}}}$ measured is --37.9° and the amplitude ratio of $\mathcal{B}_{\,C_{0.9}R_{\text{2k}}}(t)$ and $\mathcal{B}_{\,opop}(t)$ is 0.51.
Figs. 11 and 12 show the phases and amplitudes of the normalized reflected waves generated by the PZT stack under different load conditions. These figures include results from three tank tests. In each test, the position of the reflector was altered while maintaining other experimental settings constant to evaluate the consistency of the reflected signal. In both figures, the theoretical results are represented by a straight line, computed using the methods described in (18) with different load impedances at each layer, while the experimental values are obtained via the approach introduced in (14).
@@
$\frac{\mathcal{B}^{*}_{\,C_{0.9}R_{2\text{k}}}(t)}{\mathcal{B}_{\,opop}(t)}=0.45\,e^{-j\frac{\pi}{2}}\!+\!\displaystyle\frac{2000-1000}{2\left(2000+1000\right)}.$ (19)
According to (19), the phase and amplitude of the normalized $\mathcal{B}^{*}_{\,C_{0.9}R_{2\text{k}}}$ are --69.7° and 0.48, respectively. However, as illustrated in Figs. 11 and 12, the average phase of reflected waves measured through three experiments is --25.9°, and the average amplitude is 0.56.
@@
Moreover, when the loads of the two layers on the PZT stack are L0.9 and short-circuit (marked as $\text{L}_{\text{0.9}}$h in the figure), the theoretical phase and amplitude of the reflected wave are 138° and 0.67, respectively. By contrast, the average phase of reflected waves obtained from the tank tests is 102.4°, and the average amplitude is 0.71.
@@
VI-B Performance of ML-ARIS
VI-B1 Simulation settings
The simulation is conducted on the COMSOL multiphysics simulator. The reflector comprises 4 PZT-4 layers; only the first two layers are used for IQ modulation, while the remaining two remain in an open-circuit condition. Each PZT disk measures $\phi$ 60 mm $\times$ 4 mm. A monochromatic plane wave with a sound pressure of 1 Pa at 41.1 kHz serves as the signal source. Both the reflector and the water medium are surrounded by a perfect matching layer (PML) to mitigate multipath effects.
To enhance the accuracy, the simulation is structured into the following five steps: {adjustwidth}-0.71cm0cm
@@ VI-B1 Simulation settings
VI-B2 Simulation results of single reflector
Figs. 13 and 14 display the phases and amplitudes of the normalized reflected waves obtained from the simulation under various load conditions, respectively. Three measurement probes scattered in the tank are deployed to capture the reflected signal.
@@ VI-B2 Simulation results of single reflector
A comparison of Figs. 13 and 14 with Figs. 11 and 12 shows that both the phase and amplitude measured in the simulation correspond more closely with the theoretical results than those from the tank experiments. This improved accuracy is attributed to the use of PML that minimizes multipath effects, thus approximating an anechoic pool environment. Additionally, the five steps in the simulation effectively isolate the incident waves from the reflected waves, eliminating interference from direct waves on the received signal.
@@
VI-B3 Simulation results of reflector array
@@
Figure 15: Reflected beams from an ML-ARIS of 8 reflectors.
In the simulation, the ML-ARIS is configured to generate directional gain at 45° in the water through the IQ modulation. The optional status of the load network is the same as the circuit used in the tank experiments. As shown in the figure, the reflected beam exhibits significant directivity in the direction around 35°. In addition, a grating lobe near 155° can be observed as the distance between reflectors in ML-ARIS exceeds one wavelength.
Note that the absolute pressure of the reflected wave shown in Fig. 15 is small because only a small fraction of the acoustic energy propagates into the PZT due to the acoustic impedance mismatch between the water and the reflector’s head mass. In practical applications, this issue can be addressed by incorporating multiple matching layers between the head mass and the water as those used in medical [26] or sonar imaging [27] to reduce the significant reflection at the bare boundary.
@@
R1 R2 R3 R4 R5 R6 R7 R8
@@
IQ L2 1k L06 C10 L10 C10 L03 1k C06
1-bit L1 Sh Sh Op Sh Op Sh Sh Op
2-bit L1 Op Sh C10 L10 C10 Op Sh C10
@@
In Fig. 16, we compare the reflected beams generated by the ML-ARIS under various coding schemes. The configuration is identical to that in Fig. 15, and the beam amplitude is normalized to the maximum pressure collected by the probes. Each coding scheme is designed to produce a main lobe at 45°, with the corresponding load configurations provided in Table I. For the 1-bit and 2-bit coding schemes, orthogonal synthesis of the reflected signals is not required; hence, only the first layer of the PZT stack is connected to the load network, while the remaining layers are left open-circuited.
@@
From Fig. 16, it can be observed that due to near-field effects, the main lobes of all three beams are deflected to 30°, yielding a 15° deviation from the theoretical expectation. Moreover, the side lobe of reflected beam with IQ modulation is substantially lower than that of the 1-bit and 2-bit coding schemes. For example, the normalized amplitude of the first side lobe under IQ modulation is only 0.26 at 47.5°, compared with 0.52 at 45° for 2-bit coding and 0.74 at 40° for 1-bit coding. Furthermore, the sound pressure of the main lobe under IQ modulation remains strong, with normalized amplitudes of 1 compared to 0.97, and 0.86 for 1-bit, and 2-bit coding schemes, respectively.
VII Conclusion

### v2

@@ VI-B1 Experimental settings
The reflectors consist of two PZT-4 disks separated by a polyethylene terephthalate (PET) film. The sound wave propagation speed in the disk is approximately 3900 m/s; each PZT disk measures  49.2 mm diameter $\times$ 2.7 mm thickness and resonates near 44 kHz. The matching network and load states of the reflectors are controlled by the circuit board shown in Fig. 4. The conducted experiments focus on the multilayer PZT stack and its load-dependent electromechanical response, rather than a full Tonpilz transducer model. Although Fig. 3 illustrates a Tonpilz-style reflector with head and tail masses for conceptual completeness, the objective here is to isolate the impedance modulation and synthetic reflection behavior enabled by independently loaded PZT layers. For this reason, the head and tail masses are not included in the test as they primarily influence radiation efficiency and resonance characteristics, while the reflection coefficient synthesis demonstrated in this section is governed by the PZT stack and its electrical loading.
VI-B2 Preprocessing
@@
[…]
where $r_{opop}(t)$ and $r_{shsh}(t)$ represent the received signals when the two PZT disks are in open-circuit and short-circuit status, respectively. The term $s(t)$ is the original transmission signal and $\mathfrak{A}(t)\!\!=\!\!\!\sum^{N}_{i=0}\!A_{i}e^{j\psi_{i}}s(t\!-\!\tau_{i})$, where $A_{i}$, $\psi_{i}$, and $\tau_{i}$ denote the amplitude, phase, and propagation delay of the channel coefficients associated with the direct wave (i=0i\!=\!0) and the $i$th multipath signal (i>0i\!>\!0) that do not interact with the reflector. The parameters $B_{1}^{op}$ and $\phi_{1}^{op}$ specify the amplitude and phase, respectively, of the wave reflected by the first PZT layer under open-circuit status, while $B_{2}^{op}$ and $\phi_{2}^{op}$ represent those for the second PZT layer.
@@
$\mathfrak{A}(t)=\frac{r_{opop}(t)+r_{shsh}(t)}{2}.$ (14)
From (13), it can be observed that $\mathfrak{A}(t)$ does not depend on the reflector’s load status in a stable acoustic environment. Therefore, the waves reflected by the reflector for any load status can be extracted from the received signal using two reference signals. For instance, if the loads of the first and second layers of the PZT stack are $\text{C}_{\text{0.6}}$ and $\text{R}_{\text{L}}$, respectively, the received signal is denoted by $r_{C_{0.6}R_{L}}(t)$, while the signal reflected by the PZT stack is given by
$\mathfrak{B}_{\,C_{0.6}R_{L}}(t)\!=\!\displaystyle\sum^{M}_{k=1}\!\left(B^{C_{0.6}}_{1}e^{j\phi^{C_{0.6}}_{1}}\!+B^{R_{L}}_{2}e^{j\phi^{R_{L}}_{2}}\right)\!s(t-\tau_{k}).$ (15)
Based on (14), $\mathfrak{B}_{\,C_{0.6}R_{L}}(t)$ can be calculated as follows:
$\begin{array}[]{lll}\mathfrak{B}_{\,C_{0.6}R_{L}}(t)\!\!\!\!&=&\!\!\!r_{C_{0.6}R_{L}}(t)-\mathfrak{A}(t)\\ \!\!\!\!&=&\!\!\!r_{C_{0.6}R_{L}}(t)-\displaystyle\frac{r_{opop}(t)+r_{shsh}(t)}{2}.\end{array}$ (16)
The expression in (16) is used to extract the signal reflected by the reflector in a multipath environment from experimental measurements. For performance evaluation, the theoretical result of the reflected wave, $\mathfrak{B}^{*}_{\,C_{0.6}R_{L}}(t)$, is also needed. According to the circuit design presented in Fig. 4, under ideal conditions,
@@
[…]
$\mathfrak{B}^{*}_{\,C_{0.6}R_{L}}(t)\!=\!\left[0.6e^{-j\frac{\pi}{2}}\!+\!\left(\displaystyle\frac{R_{L}-Z_{0}}{R_{L}+Z_{0}}\right)\right]\displaystyle\sum^{M}_{k=1}\!s(t\!-\!\tau_{k}).$ (18)
A comparison of $\mathfrak{B}_{\,C_{0.6}R_{L}}(t)$ with $\mathfrak{B}^{*}_{\,C_{0.6}R_{L}}(t)$ enables evaluating the accuracy of the amplitude and phase of the reflected wave. For convenience, we first normalize $\mathfrak{B}_{\,C_{0.6}R_{L}}(t)$ and $\mathfrak{B}^{*}_{\,C_{0.6}R_{L}}(t)$ through $\mathfrak{B}_{\,opop}(t)$. Here, $\mathfrak{B}_{\,opop}(t)$ represents the signal reflected by the PZT stack when both layers are in open-circuit status, i.e., $\mathfrak{B}_{\,opop}(t)\!=\!\sum^{M}_{k=1}\!\left(B^{op}_{1}e^{j\phi^{op}_{1}}\!\!\!+\!B^{op}_{2}e^{j\phi^{op}_{2}}\right)\!s(t-\tau_{k})$, which can be obtained experimentally as follows:
$\mathfrak{B}_{\,opop}(t)=\displaystyle\frac{r_{opop}(t)-r_{shsh}(t)}{2}=2\displaystyle\sum^{M}_{k=1}\!s(t-\tau_{k}).$ (19)
According to the (18) and (19), the normalized $\mathfrak{B}^{*}_{\,C_{0.6}R_{L}}(t)$ can be expressed as follows:
$\frac{\mathfrak{B}^{*}_{\,C_{0.6}R_{L}}(t)}{\mathfrak{B}_{\,opop}(t)}=0.3\,e^{-j\frac{\pi}{2}}\!+\!\displaystyle\frac{R_{L}-Z_{0}}{2\left(R_{L}+Z_{0}\right)}$ (20)
Finally, when the load status of the two layers are $\text{C}_{\text{0.6}}$ and $\text{R}_{\text{L}}$, the accuracy of the reflected signal generated by the multilayer reflector in a multipath environment can be estimated by the difference between $\mathfrak{B}^{*}_{\,C_{0.6}R_{L}}(t)/\mathfrak{B}_{\,opop}(t)$, calculated by (20), with $\mathfrak{B}_{\,C_{0.6}R_{L}}(t)/\mathfrak{B}_{\,opop}(t)$, obtained experimentally via (16). This methodology can be applied to analyze the amplitude and phase of the reflection signals for arbitrary load status in the multilayered acoustic reflector.
VI-B3 Experimental results
@@
Fig. 13 (b) shows the waveform of $\mathfrak{B}(t)$, which includes the direct wave and the multiple signals that do not interact with the reflector. As described in (14), $\mathfrak{A}(t)$ is computed as half the sum of the two reference waves shown in Fig. 13 (a). Fig. 13 (c) displays the waves reflected by the PZT stack under two different load conditions, where $\mathfrak{B}_{\,C_{0.9}R_{\text{2k}}}(t)$ and $\mathfrak{B}_{\,opop}(t)$ are obtained from the two reference waveforms through the expressions introduced in (16) and (19), respectively.
To demonstrate the difference between the two reflected signals, Fig. 13 (d) provides a partially enlarged view of the area within the yellow dotted frame in Fig. 13 (c). In this figure, let $\Delta\phi_{\,C_{0.9}R_{\text{2k}}}$ denote the phase difference between $\mathfrak{B}_{\,opop}(t)$ and $\mathfrak{B}_{\,C_{0.9}R_{\text{2k}}}(t)$. Experimental results indicate that the average $\Delta\phi_{\,C_{0.9}R_{\text{2k}}}$ measured is -37.9 and the amplitude ratio of $\mathfrak{B}_{\,C_{0.9}R_{\text{2k}}}(t)$ and $\mathfrak{B}_{\,opop}(t)$ is 0.51.
Figs. 14 and 15 show the phases and amplitudes of the normalized reflected waves generated by the PZT stack under different load conditions. These figures include results from three tank tests. In each test, the position of the reflector was altered while maintaining other experimental settings constant to evaluate the consistency of the reflected signal. In both figures, the theoretical results are represented by a straight line, computed using the methods described in (20) with different load impedances at each layer, while the experimental values are obtained via the approach introduced in (16).
@@
$\frac{\mathfrak{B}^{*}_{\,C_{0.9}R_{2\text{k}}}(t)}{\mathfrak{B}_{\,opop}(t)}=0.45\,e^{-j\frac{\pi}{2}}\!+\!\displaystyle\frac{2000-1000}{2\left(2000+1000\right)}.$ (21)
According to (21), the phase and amplitude of the normalized $\mathfrak{B}^{*}_{\,C_{0.9}R_{2\text{k}}}$ are -69.7 and 0.48, respectively. However, as illustrated in Figs. 14 and 15, the average phase of reflected waves measured through three experiments is -25.9, and the average amplitude is 0.56.
@@
Moreover, when the loads of the two layers on the PZT stack are $\text{L}_{\text{0.9}}$ and short-circuit (marked as $\text{L}_{\text{0.9}}$Sh in the figure), the theoretical phase and amplitude of the reflected wave are 138 and 0.67, respectively. By contrast, the average phase of reflected waves obtained from the tank tests is 102.4, and the average amplitude is 0.71.
@@
Actually, the coupling requirements in ML-ARIS are inherently contradictory. On one hand, we seek near-perfect coupling between layers so that downstream layers receive as much incident energy as possible and can efficiently transmit their reflected waves through the upstream layers. On the other hand, we require the layers to be mechanically independent so that reflected waves with different even opposing phases do not couple through shared vibrations. Such inter-layer coupling or cross-coupling would distort synthesis.
The experimental results presented in Fig. 14 and Fig. 15 can be used to estimate the level of cross-coupling. The larger the amplitude and phase deviations from their ideal values, the greater the distortion in the synthesized signal caused by interlayer coupling. For example, when the load of two layers are identical (e.g., $\text{R}_{\text{4k4k}}$ and OpOp), the reflected wave will have the same phase and amplitude. In this case, the measured reflected wave is very close to the ideal case. However, when interlayer vibrations develop phase differences (e.g., $\text{L}_{0.9}\text{R}_{\text{2k}}$ and $\text{C}_{0.9}\text{R}_{\text{2k}}$), mechanical coupling causes the layers to interact, leading the synthesized signal to depart from the ideal response in both amplitude and phase.
Inter-layer coupling in multilayer acoustic reflectors is inherently influenced by a combination of environmental, material, and implementation-dependent factors. In underwater operation, the coupling strength depends not only on the intrinsic mechanical interaction between adjacent PZT layers, but also on external conditions such as water pressure, temperature, surrounding acoustic loading, and even the angle of the incident wave, all of which modify the effective boundary conditions and impedance of the structure. In addition, coupling is frequency-dependent and typically increases near mechanical resonance. Practical realizations further introduce variability through the choice of reflector materials, bonding layers, prestress, and assembly tolerances, as well as manufacturing-induced deviations in PZT properties. As a result, the coupling coefficient should be interpreted as an effective, environment- and configuration-specific metric rather than a universal constant, and its precise value may vary across deployments and operating conditions.
In real implementation, the inter-layer coupling issue can be addressed through calibration. By characterizing the interlayer coupling under representative operating conditions and measuring the resulting amplitude and phase deviations for each load configuration, a compensation map can be built that adjusts the programmed loads to recover the intended in-phase and quadrature components. In effect, calibration shifts the design target from the idealized independent-layer model to the true coupled response, enabling accurate signal synthesis despite residual mechanical interactions.
VI-C Performance of ML-ARIS
VI-C1 Simulation settings
The simulation is conducted on the COMSOL multiphysics simulator. The reflector comprises 4 PZT-4 layers; only the first two layers are used for synthetic reflection, while the remaining two remain in an open-circuit condition. Each PZT disk measures  60 mm diameter $\times$ 4 mm thickness. A monochromatic plane wave with a sound pressure of 1 Pa at 41.1 kHz serves as the signal source. Both the reflector and the water medium are surrounded by a perfect matching layer (PML) to mitigate multipath effects.
To enhance the accuracy, the simulation is structured into the following five steps:
@@ VI-C1 Simulation settings
VI-C2 Simulation results of single reflector
Figs. 16 and 17 display the phases and amplitudes of the normalized reflected waves obtained from the simulation under various load conditions, respectively. Three measurement probes scattered in the tank are deployed to capture the reflected signal.
@@ VI-C2 Simulation results of single reflector
A comparison of Figs. 16 and 17 with Figs. 14 and 15 shows that both the phase and amplitude measured in the simulation correspond more closely with the theoretical results than those from the tank experiments. This improved accuracy is attributed to the use of PML that minimizes multipath effects, thus approximating an anechoic pool environment. Additionally, the five steps in the simulation effectively isolate the incident waves from the reflected waves, eliminating interference from direct waves on the received signal.
@@
VI-C3 Simulation results of reflector array
@@
Figure 18: Reflected beams from an ML-ARIS with 8 reflectors.
@@
Reflectors 1 2 3 4 5 6 7 8
Layer 1 Open $\text{R}_{0.8\text{k}\Omega}$ $\text{R}_{2.8\text{k}\Omega}$ $\text{R}_{1.1\text{k}\Omega}$ $\text{R}_{0.28\text{k}\Omega}$ $\text{R}_{19.4\Omega}$ Short $\text{R}_{9.4\text{k}\Omega}$
IQ Layer 2 $\text{R}_{1\text{k}\Omega}*$ $\text{L}_{\text{0.6}}$ $\text{C}_{\text{1.0}}$ $\text{L}_{\text{1.0}}$ $\text{C}_{\text{1.0}}$ $\text{L}_{\text{0.3}}$ $\text{R}_{1\text{k}\Omega}$ $\text{C}_{\text{0.6}}$
1-bit Layer 1 Short Short Open Short Open Short Short Open
2-bit Layer 1 Open Short $\text{C}_{\text{1.0}}$ $\text{L}_{\text{1.0}}$ $\text{C}_{\text{1.0}}$ Open Short $\text{C}_{\text{1.0}}$
@@
Inductive load for a reflected-signal phase of ++90 and magnitude of 0.6, computed from (2).
Capacitive load for a reflected-signal phase of -90 and magnitude of 1.0, computed from (2).
@@
In Fig. 19, we compare the reflected beams generated by the ML-ARIS under various coding schemes. The configuration is identical to that in Fig. 18, and the beam amplitude is normalized to the maximum pressure collected by the probes. Each coding scheme is designed to produce a main lobe at 45, with the corresponding load configurations provided in Table I. For the 1-bit and 2-bit coding schemes, orthogonal synthesis of the reflected signals is not required; hence, only the first layer of the PZT stack is connected to the load network, while the remaining layers are left open-circuited.
@@
From Fig. 19, it can be observed that due to near-field effects, the main lobes of all three beams are deflected to 30, yielding a 15 deviation from the theoretical expectation. Moreover, the side lobe of reflected beam with synthetic reflection is substantially lower than that of the 1-bit and 2-bit coding schemes. For example, the normalized amplitude of the first side lobe under synthetic reflection is only 0.26 at 47.5, compared with 0.52 at 45 for 2-bit coding and 0.74 at 40 for 1-bit coding. Furthermore, the sound pressure of the main lobe under synthetic reflection remains strong, with normalized amplitudes of 1 compared to 0.97, and 0.86 for 1-bit, and 2-bit coding schemes, respectively.
Typical long-range UA-RIS deployments are expected to operate in the far field at the frequencies considered. Accordingly, links with ranges well beyond the Fraunhofer distance fall primarily in the far-field regime, whereas near-field calibration is most critical for close-in configurations (e.g., tank experiments or short-range installations) where near-field beam deviations become non-negligible.
The SNR gains achievable with ML-ARIS are expected to exceed those reported in our prior experimental UA-RIS study [13]. Both ML-ARIS and [13] synthesize reflections using two orthogonal (I/Q) components; however, as discussed in Section. III, the multilayer architecture effectively doubles the number of controllable reflecting layers per unit, yielding a larger effective aperture (or, equivalently, more degrees of freedom) for the same number of reflection units. For reference, in [13] lake trials, deploying a UA-RIS with 24 reflecting units near the transmitter extended the communication range by up to 28% and 46% in deep- and shallow-water environments, respectively. At a fixed transmitter-receiver separation, the UA-RIS increased the received SNR by 2.13 dB on average, with peak gains up to 2.92 dB in certain cases. When the UA-RIS was positioned near the receiver, the communication range increased by 40.6% and 66% in deep and shallow water, respectively; for a fixed separation, the received SNR improved by 2.56 dB on average and by as much as 4.2 dB under specific conditions.
VII Conclusion and Future Work